LINEAR EQUATIONS
A. Solve for the unknown in each equation.
4(x+5) = 6(x - 2)

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

A. Solve for the unknown in each equation.
4(x+5) = 6(x - 2)
Distribute
4x + 20 = 6x - 12
20 + 12 = 6x - 4x
32 = 2x
x = 32/2
x = 16
:
:
Check in original equation
4(16+5) = 6(16-2)
4(21) = 6(14)
84 = 84

Answer by ikleyn(52800)

.

If you need to learn on how to solve linear equations in one unknown, read the lesson
    - HOW TO solve a linear equation
in this site.